The sum of two integers is 64. One integer is 20 less than twice the other.
Find the smaller integer.

Answer by fractalier(6550)

Call the integers x and y. Then we have
x + y = 64 and
x = 2y - 20
Substitute the second into the first equation and get
2y - 20 + y = 64
3y - 20 = 64
3y = 84
y = 28
x = 36
The smaller is 28.